If sin x + cos x=1 , then prove that sin x - cos x=√1

Grade:11

2 Answers

Adithya prakash
8 years ago
i think the the question is wrong
since 10.5 = 1
sinx-cosx=1
therefore,
sinx + cosx= sinx-cosx
which is not possible
excuse me if im wrong.
 
Sumanth
7 years ago
Sin x + cos x =1Sin x=1-cos x2sin(x/2)cos(x/2)=2sin^2(x/2)Cos(x/2)=sin(x/2)1=tan(x/2)Tan45=tan(x/2)X/2=45X=90Sin (90) -cos(90)= 1-0= 11=√1Hence proved